Fermi National Laboratory
 

Accelerator Update for March 30, 2007 - April 2, 2007

 

Friday March 30

The day shift began with the Tevatron (TeV) recovering from a quench at sector F4, with Operations monitoring the Recycler stash (275.7E10), with the Antiproton Source (Pbar) stacking (30.3E10), with the Booster, Linac, Main Injector (MI), and Recycler operating normally, with MiniBooNE off, with NuMI and the Meson MTest experiments taking beam, and with D-Zero (D0) and the Collider Detector at Fermilab (CDF) waiting for the colliding beams.

Operations prepared D0 and CDF for access. They were both in their collision halls by 8:57 AM. CDF was out of access about 30 minutes later.

The Neutron Therapy Facility (NTF) reported at 10:14 AM that they had completed their treatments for the day.

Operations put the TeV through dry squeeze at 10:44 AM.

D0 came out of access at 10:59 AM.

Operations put the TeV through a 4x0 wet squeeze at 11:38 AM.

Operations put the TeV into a shot setup at 12:43 PM.

At 2:34 PM, the TeV suffered a quench. A sector F3 bulk power supply had tripped. When an operator turned the supply on, the TeV quenched at sectors A4, B1, C4, D1, F4, and all the low beta magnets.

Operations reported at 4 PM that the Recycler had stashed antiprotons twice during the day shift.

The MTest experimenters reported at 6:12 PM that they were through with the beam for the night.

Operations began the TeV turn on at 6:51 PM.

Operations put the TeV through a dry squeeze at 7:39 PM.

TeV experts began a sextupole study at 8:25 PM.

Operations reported at 11:58 PM that the Recycler had stashed four times during the evening shift.

Saturday March 31

The midnight shift began with the TeV conducting a study, with Operations monitoring the stash (196.9E10), with Pbar stacking (11.6E10), and with NuMI taking beam.

Operations put the TeV through a wet squeeze at 1:15 AM.

TeV studies ended around 1:30 AM.

Operations put the TeV into shot setup at 2:25 AM.

Operations established store 5314 at 4:19 AM with an initial luminosity of 196.4E30.

Operations established beam to MTest at 6:01 AM.

Operations reported at 7:58 AM that the Recycler had stashed once during the midnight shift.

At 8:35 AM, Operations reported that an Accumulator kicker (A:IKIK) stopped pulsing. Operations called in an expert.

At 9 AM, the H- Source vacuum gauge tripped off. Operators accessed the H- dome and reset the device.

The kicker expert reported at 10:03 AM that the kicker was shorted and the short was on a cable coming out of the power supply. He had it repaired by 11:24 AM. Stacking resumed.

MTest reported that they were done with beam for the night at 6:03 PM.

Operations reported at 8:05 PM that the Recycler had stashed four times during the last 12 hours. Status report: the store luminosity was 49.34E30, the stack was at 25.8E10, and the stash was at 198.8E10.

At 9:30 PM, a Booster RF station (BRF7) tripped off. It reset.

At 9:32 PM, a Booster kicker (B:MKS7) tripped off and at first it wouldn’t reset. When operators did get it on they could hear sparking. An expert came in and at 11:14 PM reported that he had repaired a bad PFN cable.

Sunday April 1

The midnight shift began with Operations monitoring store 5314 and the stash, with Pbar stacking, and with NuMI taking beam.

At 12:36 AM, a Linac RF station (LRF3) tripped off for the first of three times in about thirty minutes. It reset.

Operations terminated store 5314 at 2:01 AM.

Operations put the TeV into shot setup at 2:14 AM.

Operations established store 5315 at 4:38 AM with an initial luminosity of 211.14E30.

Operations established beam to MTest at 6:56 AM.

Operations reported at 7:55 AM that the Recycler had stashed three times during the last 12 hours. Status report: the store luminosity was 126.32E30, the stack was at 48.7E10, and the stash was at 49.1E10.

At 8:49 AM, the H- Source tripped off. It reset.

At 11:27 AM, the MI beam permit tripped off due to a bad beam position for the first of three times in about two hours. It reset.

At 2:12 PM, MI tripped off due to bad beam positions for the first of many times over the next four hours. It reset each time.

At 2:47 PM, LRF3 tripped off. Operators investigated, raised the screen voltage, and had the station back on by 3:04 PM.

At 4:03 PM, LRF4 tripped off. It reset.

At 4:16 PM, store 5314 quenched at multiple locations. The fault was a ground fault on the electrical safety system. Interlock technicians came it to investigate.

At 5:56 PM, Operations reported that a Pbar Debuncher kicker (D:EKIK) wasn’t always firing with the other kickers. Operations called in experts.

At 6:29 PM, D:EKIK stopped firing. Operations loaded a NuMI only time line into the time line generator (TLG) while the experts investigated.

Operations reported at 8 PM that the Recycler had stashed three times during the last 12 hours. Status report: the stack was at 29.4E10, and the stash was at 198.9E10.

Pbar resumed stacking at 8:10 PM. The experts reported that they replaced a firing module for D:EKIK.

At 10:59 PM, cryo system techs reported that their nitrogen re-liquefying plant at CHL went down.

At 11:25 PM, operators accessed the H- dome to reset its vacuum gauge. Beam resumed about 45 minutes later.

The interlock experts reported that they traced the TeV safety system fault to TeV sector A48, and then the fault went away.

Monday April 2

The midnight shift began with the TeV recovering from a quench, with Operations monitoring the stash, with Pbar stacking, and with NuMI taking beam.

At 12:52 AM, a MI vacuum pump (M:IP107) tripped off. Operations reset it and got it back on about 20 minutes later.

At 1:12 AM, Operations reported that M:IP107 had tripped off for the fourth time. Operations left the vacuum pump off.

At 1:55 AM, Operations reported that they were having problems closing the VCBs. The Duty Electrician reported that the switch was open at the master substation and there were three trip flags. Operations called the High Voltage Coordinator.

At 3:55 AM, Operations reported that when they closed the VCB, the TeV sector B3 bus-work cabinet faulted. Operations bypassed the B3 power supply and experts reconfigured the TeV supplies to run without B3.

Operations established beam to MTest at 6:21 AM.

Operations began ramping the TeV at 6:45 AM.

Operations reported at 8:03 AM that the Recycler had stashed four times during the last 12 hours.

The Plans for Monday

The plans for the morning are to establish a store and keep it running until 6:30 AM on Tuesday. At 7 AM on Tuesday (4/3/07) the TeV will begin a 10-12 hour shutdown so that CDF can drain and refill their cooling system. The Run Coordinator will allow other machine work that doesn’t interfere too much with stacking and stashing.


Accelerator Update Archive

More Information
For Tevatron luminosity charts and the current status of Fermilab's accelerators and detectors (live!), please go to Fermilab Now.

Comments and Suggestions
What do you think about the Accelerator Updates? Please send comments and suggestions to: accelupdates@fnal.gov.

last modified 04/02/2007    email Fermilab